### v1

# Irreducible finite-index representations of automorphism groups of trees

Abstract
We consider a locally finite tree $T$ where every vertex has degree 3$\geq$ 3≥ 3, and a closed noncompact subgroup $G$ of the automorphism group $\mathrm{Aut}(T)$, with Tits’ independence property (P) and with transitive action on the boundary $\partial T$. We show that $G$ does not have continuous irreducible representations to a locally convex topological vector space over $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{H}$ that preserve a continuous nondegenerate sesquilinear form of index $p>1$. As the representations of index 111 are already classified, this gives a complete classification of continuous irreducible representations of $G$ that preserve a continuous nondegenerate sesquilinear form of finite index.
1 Introduction
In this paper, we study automorphism groups of locally finite trees and their continuous irreducible linear representations. We focus on a class of groups that satisfy Tits’ independence property (P) ([19], 4.2) and act transitively on the boundary at infinity of the tree. We assume that every vertex of the tree has degree 3$\geq$ 3≥ 3, in which case the hypotheses imply that the tree is homogeneous or semi-homogeneous. The groups we consider include universal groups associated with 222-transitive permutation groups, which were introduced by Burger and Mozes in ([4], 3.2) and then studied by Amann in his Ph.D. thesis ([1], Chapter 3). More recently, Reid and Smith [18] provided a characterization of tree automorphism groups with Tits’ independence property (P) in terms of combinatorial structures called local action diagrams.
In his Ph.D. thesis ([1], Chapter 2), Amann has classified continuous irreducible unitary representations of these groups, which fall into three families: super-cuspidal, special and spherical, according to the conjugacy class of a maximal compact subgroup that fixes a nontrivial vector. Our goal here is to classify continuous irreducible representations that are not unitary but preserve a continuous nondegenerate sesquilinear form of finite index instead. From now on, all sesquilinear from will be implicitly assumed to be continuous.
In general, we will work on a locally compact topological vector space over the field of real, complex, or quaternionic numbers. The existence of a sesquilinear form implies that the space has a pre-Hilbert structure, but we do not assume completeness. Irreducible representations are meant in the algebraic sense: the representation does not preserve any proper vector subspace except the trivial one. Our representations can also be regarded as continuous homomorphisms of $G<\mathrm{Aut}(T)$, where $T$ is a homogeneous or semi-homogeneous tree and $G$ satisfies the above assumptions, to $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, or $\mathrm{Sp}(p,\infty)$, with $p\geq 1$.
The study of Hilbert spaces with indefinite bilinear forms dates back to Pontrjagin [17], Iohvidov-Krein [10], and Naimark [13, 15, 12, 14]. More recently, Gromov ([9], Section 6) has suggested the study of the infinite-dimensional symmetric spaces of finite rank that can be defined subsequently. The case of the infinite-dimensional real hyperbolic space has been studied by Burger, Iozzi and Monod in [3], and later by Monod and Py in [11], while the general case has been studied by Duchesne in recent times [5, 8, 7, 6]. It turns out ([7], Corollary 1.10) that the only irreducible infinite-dimensional symmetric spaces of non-positive curvature and finite rank are, up to homothety, $\mathrm{O}(p,\infty)/\mathrm{O}(p)\times\mathrm{O}(\infty)$, $\mathrm{U}(p,\infty)/\mathrm{U}(p)\times\mathrm{U}(\infty)$, and $\mathrm{Sp}(p,\infty)/\mathrm{Sp}(p)\times\mathrm{Sp}(\infty)$, and that their isometry groups are, up to a subgroup of index two in the complex case, exactly the projective versions of $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, and $\mathrm{Sp}(p,\infty)$ ([6], Theorem 3.3). Therefore, studying representations of $G$ that preserve a nondegenerate sesquilinear form of finite index is the same as studying actions of $G$ by isometries on infinite-dimensional symmetric spaces of nonpositive curvature and finite rank.
In another paper [20], we consider the case where $G$ is a higher-rank algebraic group over a non-archimedean local field $\mathbb{F}$, and we show that, at least in the case $G=\mathrm{SL}_{n}(\mathbb{F})$ ($n\geq 3$), there are no continuous fixed point-free actions of $G$ on any infinite-dimensional symmetric space of nonpositive curvature and finite rank, which implies that there are no continuous irreducible representations of $G$ that preserve a nondegenerate sesquilinear form of finite index.
Turning back to the case where $G$ is a subgroup of the automorphism group of a tree, continuous irreducible representations to $\mathrm{O}(1,\infty)$ correspond to continuous actions by isometries on the infinite-dimensional real hyperbolic space, and these have been classified by Burger, Iozzi and Monod ([3], Theorem C), at least in the case where the tree $T$ is homogeneous and the action of $G$ is triply transitive on the boundary $\partial T$. The classification consists of a family of representations that depend on a single real parameter.
Our main result in this paper is that the classification by Burger, Iozzi and Monod basically exhausts all possible continuous irreducible representations of finite index. In fact we will show that, if $p>1$, there exist no continuous irreducible representations of $G$ to $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, or $\mathrm{Sp}(p,\infty)$, and that in the case $p=1$ all complex and quaternionic representations are obtained by taking the tensor product with $\mathbb{C}$ or $\mathbb{H}$ of a real representation. The proof is done by generalizing the classification carried out by Amann ([1], Chapter 2) to the case where the representations are not necessarily unitary. Our classification is a generalization of the one carried out by Ol’shanskii [16] for $G=\mathrm{Aut}(T)$, except that we do not dive into the details of unitary representations but focus on those that can preserve indefinite sesquilinear forms.
The paper is organized as follows. In Section 2 we give a few preliminaries on trees and on the independence property, which allow us to precisely state our hypotheses and the thesis of our theorem. In Section 3 we present the classification of the representations as spherical, special, and super-cuspidal, and we show that for our purposes it suffices to consider spherical representations. In Section 4 we consider the case where $G$ acts transitively on the set of vertices of the tree and give an explicit description of spherical representations, getting a classification that depends on only one parameter. In Section 5 we study nondegenerate sesquilinear forms of finite index that can be preserved by these representations, completing the proof of our theorem in the case where the action of $G$ is transitive on the vertices. Finally, in Section 6 we generalize the results of Sections 4 and 5 to the case where the action of $G$ is not transitive on the vertices.
This paper is part of my Ph.D. project under the supervision of Prof. Nicolas Monod at EPFL (Lausanne, Switzerland). I am grateful to Nicolas for proposing this project, for giving me valuable advice and feedback, and for sharing with me deep insights of the theory. I am also grateful to Bruno Duchesne for his useful questions and remarks which helped me make this paper more clear.
2 Preliminaries
[…]
The boundary at infinity $\partial T$ of the tree $T$ is the set of infinite injective paths $(x_{n})_{n\in\mathbb{N}}$ with $x_{n}\in V$ for $n\in\mathbb{N}$, where two paths $(x_{n})_{n\in\mathbb{N}}$ and $(y_{n})_{n\in\mathbb{N}}$ are identified if there exists $t\in\mathbb{Z}$ such that $x_{n}=y_{n+t}$ for all large enough $n$.
We consider a locally finite tree $T$ where every vertex has degree 3$\geq$ 3≥ 3, and its automorphism group $\mathrm{Aut}(T)$. This group, equipped with the topology of pointwise convergence, is totally disconnected and locally compact. It has a natural action on the tree and on its boundary at infinity.
Let $G<\mathrm{Aut}(T)$ be a subgroup. For each subtree $S\subseteq T$ (which might be a single vertex or an edge), let $G(S)$ be the subgroup of $G$ that fixes every point in $S$.
If $e\in E$ is an edge, it divides the tree into two subtrees whose intersection is just $e$ (together with its two endpoints). We call the two subtrees $T_{e}^{(1)}$ and $T_{e}^{(2)}$. It is immediate to see that every automorphism of $T$ that fixes every point in $T_{e}^{(1)}$ commutes with every automorphism of $T$ that fixes every point in $T_{e}^{(2)}$.
We are ready to introduce the independence property for a subgroup $G<\mathrm{Aut}(T)$. We use the definition given by Amann ([1], Definition 9):
Definition 2.1.
(Independence property) Let $G<\mathrm{Aut}(T)$ be a subgroup. We say that $G$ has the independence property if, for every edge $e$, it holds
$G(e)=G(T_{e}^{(1)})G(T_{e}^{(2)}).$
As observed by Amann (as a consequence of Lemma 10 in [1]), if $G$ is a closed subgroup of $\mathrm{Aut}(T)$, this property is equivalent to Tits’ Property (P), which was defined in ([19], 4.2).
We are interested in studying representations of closed noncompact subgroups of $\mathrm{Aut(T)}$ that act transitively on the boundary $\partial T$ and have the independence property. From now on, let $G$ be such a subgroup. We recall a few properties of $G$ that follow from the assumptions; more details can be found in [1] (Proposition 4).
Proposition 2.2.
The following hold:
  1. (a)
    The identity in the group $G$ has a basis of open compact neighborhoods: the subgroups $(G(B_{n}(o)))_{n\in\mathbb{N}}$, where $o$ is a fixed vertex and $B_{n}(o)$ is the ball of radius $n$ around $o$.
  2. (b)
    For every vertex $o$, the subgroup $G(o)$ acts transitively on $\partial T$.
  3. (c)
    Either the action of $G$ on $V$ is transitive or it has two orbits. In the first case, the tree is homogeneous; in the second case, the tree is homogeneous or semi-homogeneous, one orbit is the set of vertices at even distance from a fixed vertex $o$, and the other orbit is the set of vertices at odd distance from $o$.
We are ready to state our main theorem. We use the notation $(\pi,H)$ for representations of $G$, where $H$ is a locally convex topological vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$, and $\pi:G\rightarrow\mathrm{Aut}(H)$ is a group homomorphism. To avoid ambiguity, in the case $\mathbb{K}=\mathbb{H}$ we say that we consider a right vector space (i. e. where the scalar multiplication is on the right), so that actions are on the left.
Theorem 2.3.
Let $T$ be a locally finite tree where every vertex has degree 3$\geq$ 3≥ 3. Let $G$ be a closed noncompact subgroup of $\mathrm{Aut}(T)$, with the independence property and transitive action on $\partial T$. Let $(\pi,H)$ be a continuous irreducible $\mathbb{K}$-representation of $G$ to a locally convex topological vector space. Suppose that there exists a continuous nondegenerate finite-index sesquilinear form $B$ on $H$ that is preserved by $\pi$. Then the index of $B$ is 0 or 111. Moreover, if the index is 111, then the representation $(\pi,H)$ is isomorphic to the tensor product with $\mathbb{K}$ of a real representation.
We start with a simple but important observation.
Proposition 2.4.
Let $(\pi,H)$ be a continuous irreducible representation of $G$. Then $(\pi,H)$ is smooth, i. e. every $v\in H$ is fixed by some open compact subgroup of $G$.
Proof.
Since the representation is irreducible and the subspace of smooth vectors, i. e. vectors that are fixed by some open compact subgroup of $G$, is $G$-invariant (if $g\in G$ and $K<G$ satisfies $Kv=v$, then $gv$ is fixed by the subgroup $gKg^{-1}$), it is enough to show that there exists one nonzero smooth vector.
[…]
3 Restricting to spherical representations
In this section, we begin the classification of the representations. First, we define:
Definition 3.1.
(Complete subtree) A finite subtree $S\subset T$ is complete if every vertex in $S$ with at least two neighbors in $S$ is such that all its neighbors in $T$ are in $S$.
Given a continuous representation $(\pi,H)$ of $G$, we know from Proposition 2.4 that there is some open compact subgroup $K<G$ such that $H^{K}\neq\{0\}$, where the notation indicates the subspace of $K$-fixed vectors in $H$. We can consider the set of complete subtrees $S$ such that $H^{G(S)}\neq\{0\}$. What we have said, together with Proposition 2.2(a), implies that this set is not empty. Thus, we can consider the set $M$ of minimal complete subtrees $S$ in the set and distinguish three cases:
  • $M$ contains a subtree that consists of only one vertex: we say that the representation $(\pi,H)$ is spherical.
  • $M$ contains a subtree that consists of two vertices (linked by an edge): we say that the representation $(\pi,H)$ is special.
  • $M$ contains a subtree with more than two vertices: we say that the representation $(\pi,H)$ is super-cuspidal.
Since the set $M$ is $G$-invariant, it is clear that the representation $(\pi,H)$ belongs to exactly one of the three cases (it follows from (b) and (c) of Proposition 2.2).
We show that every irreducible continuous representation of $G$ is a subrepresentation of a more concrete representation. This will allow us to restrict our analysis to spherical representations and to express them in a more convenient way.
Proposition 3.2.
Let $S$ be a complete finite subtree of $T$. Let $C(S)$ be the space of $\mathbb{K}$-valued functions on $G$ that are right-invariant under $G(S)$ and left-invariant under some open compact subgroup of $G$ (which may depend on the function). Consider the representation $(\Pi_{S},C(S))$ of $G$ defined by the natural left action of $G$ on $C(S)$. Let $(\pi,H)$ be a continuous irreducible representation of $G$ such that $H^{G(S)}\neq\{0\}$. Then $(\pi,H)$ is a subrepresentation of $(\Pi_{S},C(S))$.
Proof.
Consider the contragredient representation $(\widetilde{\pi},\widetilde{H})$ of $(\pi,H)$. Since $H^{G(S)}\neq\{0\}$, we have $\widetilde{H}^{G(S)}\neq\{0\}$: the space $H^{G(S)}$ has a $G(S)$-invariant complement in $H$ and hence is isomorphic to $\widetilde{H}^{G(S)}$ ([2], Proposition 2.7).
Consider a nonzero $\xi\in\widetilde{H}^{G(S)}$. The map
$v\rightarrow\left(g\mapsto\left\langle\xi,\pi(g^{-1})v\right\rangle\right)$
is a well-defined map from $H$ to $C(S)$: the function $g\mapsto\left\langle\xi,\pi(g^{-1})v\right\rangle$ is right-invariant under $G(S)$ because $\xi$ is invariant under $G(S)$, and left-invariant under some open compact subgroup of $G$ because every $v\in H$ is smooth. It is immediate to see that the map is $G$-equivariant and nonzero. Therefore, since the representation $(\pi,H)$ is irreducible, the map is injective and $(\pi,H)$ is a subrepresentation of $(\Pi_{S},C(S))$.
Amann proves ([1], Lemma 20) that if the representation $(\pi,H)$ is super-cuspidal, the space $H\subset C(S)$ that arises by applying Proposition 3.2 to $S\in M$ consists of functions that are finitely supported when projected to $G/S$. This is shown for complex representations, but the proof is the same for real and quaternionic ones. Amann then proceeds to classify unitary super-cuspidal representations, but the proof that the functions are finitely supported on $G/S$ works without assuming that the representation is unitary. This follows as a consequence, as the representation will preserve the scalar product on $H$ defined as the restriction to $H$ of the standard scalar product on the space of finitely supported functions on $G/S$.
Similarly, Amann proves ([1], Lemma 34) that if the representation $(\pi,H)$ is special, all functions in $H\subset C(S)$ (with $S$ consisting of two vertices linked by an edge) belong to $\ell^{2}(G/S)$. Again, the proof does not need that the representation is unitary, and this fact follows as a consequence, as the restriction to $H$ of the standard scalar product on $\ell^{2}(G/S)$ will be preserved.
In the rest of the paper, we will focus our attention on the remaining case, the case of spherical representations. Since in the other two cases all irreducible representations are unitary and since they satisfy $\dim(H^{G(S)})<\infty$ ([1], Lemma 20, Lemma 34), they cannot preserve a nondegenerate sesquilinear form of index 1$\geq$ 1≥ 1. This is explained in the following proposition.
Proposition 3.3.
[…]
Proof.
[…]
$\langle A\pi(g)v,w\rangle=\langle\langle\pi(g)v,w\rangle\rangle=\langle\langle v,\pi(g^{-1})w\rangle\rangle=\langle Av,\pi(g^{-1})w\rangle=\langle\pi(g)Av,w\rangle$
for all $g\in G,\ v,w\in H$. This means that $A$ commutes with the representation, i. e. $A\in\mathrm{End}(\pi)$. By the Schur Lemma ([2], Lemma 3.21), $\mathrm{End}(\pi)$ is a division algebra over $\mathbb{R}$. If we fix a nonzero $v\in H^{K}$ the map $\mathrm{End}(\pi)\rightarrow H^{K}$ defined by $\phi\mapsto\phi(v)$ is injective, since $\pi$ is irreducible, thus $\dim(\mathrm{End}(\pi))\leq\dim(H^{K})$. Since $\dim(H^{K})$ is finite, so is $\dim(\mathrm{End}(\pi))$, and it is well known that then $\mathrm{End}(\pi)$ must be isomorphic to $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{H}$.
If $\mathrm{End}(\pi)$ is isomorphic to $\mathbb{R}$, then we are done. In the other two cases, we have $A=\alpha Id+\beta A_{1}$ with $\alpha,\beta\in\mathbb{R}$ and $A_{1}^{2}=-Id$. We wish to prove that $\beta=0$. Since $A$ and $\alpha Id$ are self-adjoint and both belong to $\mathrm{End}(\pi)$, the same must be true for $\beta A_{1}$ and therefore, if $\beta\neq 0$, also for $A_{1}$. Then for every $v\in H$ we have
$\langle A_{1}v,A_{1}v\rangle=\langle A_{1}^{2}v,v\rangle=-\langle v,v\rangle.$
[…]
4 An explicit form for spherical representations
If the representation $(\pi,H)$ is spherical, we can apply Proposition 3.2 with $S$ being a single vertex. We call this vertex $o$, and we set $K:=G(o)$.
The space $C(S)$ is the space of $K$-right-invariant functions on $G$ that are left-invariant under some open compact subgroup. We can identify this space with the space of functions on $G/K$ (the set of right cosets of $G$ under $K$) that are left-invariant under some open compact subgroup of $G$. The set $G/K$ can be identified with the orbit of $o$ in $T$ under the action of $G$, which is either the full vertex set $V$ or the set of vertices at even distance from $o$ (Proposition 2.2(c)).
In this section, we assume that we have the full vertex set $V$, i. e. that the action of $G$ is transitive on the vertices. In this case, the tree must be homogeneous. In the last section of the paper, we will extend the results to the other case.
Proposition 3.2 now tells us that we can see the representation $(\pi,H)$ as a subrepresentation of $(\Pi,C)$, where $C$ is the space of functions on $V$ that are left-invariant under some open compact subgroup of $G$ and $\Pi$ is the left action of $G$ on $C$. Therefore, from now on we assume that $H\subset C$ and $\pi$ is the restriction of $\Pi$ to $H$.
It is known that $(G,K)$ is a Gel’fand pair, which means that the convolution algebra of compactly supported complex functions on $K\backslash G/K$ is commutative. This follows ([1], Proposition 7) from the stronger fact that for every $g\in G$ we have $g^{-1}\in KgK$: the set $KgK$ is the set of $h\in G$ such that $d(ho,o)=d(go,o)$, and we have $d(g^{-1}o,o)=d(o,go)$.
Proposition 4.1.
For any irreducible representation $(\pi,H)$ of $G$ that preserves a sesquilinear form of finite index $q$, the space $H^{K}$ of $K$-fixed vectors is at most 111-dimensional.
Proof.
We consider the convolution algebra $\mathcal{A}$ of compactly supported functions on $K\backslash G/K$ (i. e. compactly supported functions on $G$ which are invariant under $K$ on the left and on the right). We consider real functions if $\mathbb{K}=\mathbb{R}$, and complex functions if $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{H}$.
To prove our proposition, we follow the proof of Proposition 5.4 in [11]. As in that proof, we observe that this algebra is commutative and symmetric, the latter meaning that the adjoint of every element of $\mathcal{A}$ (with respect to the sesquilinear form preserved by $\pi$) is again in $\mathcal{A}$. Since in our case $g^{-1}\in KgK$ for every $g\in G$, a stronger statement holds: every element of $\mathcal{A}$ is self-adjoint.
As in [11] we see that $\mathcal{A}$ acts irreducibly on $H^{K}$ (which we assume to be non-trivial), we see that the restriction of the sesquilinear form on $H$ to $H^{K}$ is non-degenerate (say of index $q^{\prime}\geq 0$), and we use the fact, due to Naimark ([13], Corollary 2) that a commutative symmetric algebra of operators of a complex Hilbert space endowed with a sesquilinear form of finite index $q^{\prime}\geq 1$ preserves a non-negative subspace $L$ of finite dimension. This is also true in the case $q^{\prime}=0$, due to the Schur Lemma, as explained in the addendum to [11].
In the case $\mathbb{K}=\mathbb{C}$, the result in [13] applied to the representation of $\mathcal{A}$ on $H^{K}$ ensures that $\mathcal{A}$ preserves a finite-dimensional subspace $L$ of $H^{K}$, and as the representation is irreducible it must be $L=H^{K}$. Since $\mathcal{A}$ is commutative every element of $\mathcal{A}$ is an endomorphism of the representation, so the Schur Lemma implies that $\mathcal{A}\cong\mathbb{C}$ and therefore $H^{K}$ is 111-dimensional.
In the case $\mathbb{K}=\mathbb{H}$, we apply the result in [13] to the complex representation of $\mathcal{A}$ on $H^{K}$ seen as a complex vector space and find that it preserves a finite-dimensional complex subspace $L$. The quaternionic vector space that it generates is finite-dimensional and must coincide with $H^{K}$ by irreducibility. By applying the Schur Lemma, we see again that $\mathcal{A}\cong\mathbb{C}$ and therefore $H^{K}$ must be 111-dimensional.
In the case $\mathbb{K}=\mathbb{R}$, we apply the result in [13] to the complexification of the representation of $\mathcal{A}$ on $H^{K}$. We find that there exists a finite-dimensional complex vector space $L$ preserved under $\mathcal{A}\otimes\mathbb{C}$ and thus under $\mathcal{A}$. This implies that there exists a finite-dimensional real vector space $L^{\prime}\subseteq H^{K}$ preserved under $\mathcal{A}$, which must coincide with $H^{K}$ by irreducibility. The representation preserves a positive definite bilinear form on $L^{\prime}=H^{K}$, since the restriction of the sesquilinear form to $L$ is non-negative and non-degenerate. If we apply the Schur Lemma, we see that $\mathcal{A}$ is a finite-dimensional commutative division algebra over $\mathbb{R}$ and therefore it is isomorphic to $\mathbb{R}$ or $\mathbb{C}$. In the latter case, there exists $A\in\mathcal{A}$ such that $A^{2}=-Id$, which implies, since each $A\in\mathcal{A}$ is self-adjoint, that for every $v\in H^{K}$ it holds
$\langle Av,Av\rangle=\langle A^{2}v,v\rangle=-\langle v,v\rangle,$
contradicting the fact that the form is positive definite. Therefore, it must be $\mathcal{A}\cong\mathbb{R}$, which implies that $H^{K}$ is 111-dimensional.
Proposition 4.1 implies that there exists a function $f\in H$ such that the space of $K$-invariant functions in $H$ is the space of scalar multiples of $f$.
Observation 4.2.
A function in $C$ is $K$-invariant if and only if it is radial with respect to the base vertex $o$, i. e. its value at a vertex $v$ only depends on the distance between $v$ and $o$. This follows from the fact that $K$ acts transitively on $\partial T$ (Proposition 2.2(b)). Therefore, we can use the notation $f(n)$ ($n\in\mathbb{N}$) to indicate the value of $f$ at any vertex at distance $n$ from $o$.
Definition 4.3.
(Laplacian operator) The Laplacian operator on the tree $T$ is the operator $L:C\rightarrow C$ defined as follows: if $f:V\rightarrow\mathbb{K}$ is a function in $C$, the function $Lf$ associates to every vertex $v$ in $V$ the average of the values of $f$ at neighbors of $v$.
Observation 4.4.
The Laplacian $Lf$ of a radial function $f$ is a radial function, with the following values ($d$ is the degree of the tree):
  • $Lf(0)=f(1)$,
  • $Lf(n)=\frac{1}{d}f(n-1)+\frac{d-1}{d}f(n+1)$ for all $n\geq 1$.
Proposition 4.5.
The function $f$ is an eigenfunction of the Laplacian, i. e. there exists some $\alpha\in\mathbb{K}$ such that $Lf=\alpha f$.
Proof.
Since the action of $G$ on $V$ is transitive, for every neighbor $w$ of $o$ there is an element $g\in G$ such that $go=w$. If we consider the function $\pi(g)f$, it is a radial function with respect to $w$: for every $n$, the value of $\pi(g)f$ at any vertex at distance $n$ from $w$ is equal to $f(n)$.
We repeat the process for all neighbors of $o$, and we average the functions $\pi(g)f$ that we obtain. We get a function $Af\in H$, radial with respect to $o$. We have:
  • $Af(0)=f(1)$,
  • $Af(n)=\frac{1}{d}f(n-1)+\frac{d-1}{d}f(n+1)$ for all $n\geq 1$.
Therefore, due to Observation 4.4, $Af=Lf$.
Since the space of radial functions in $H$ (which is the same as the space of $K$-invariant functions in $H$) is 111-dimensional, and both $f$ and $Af$ are in $H$, we must have $Af=\alpha f$, and hence $Lf=\alpha f$, for some $\alpha\in\mathbb{K}$.
It is immediate to see, from Observation 4.4, that $f$ is uniquely determined (up to a scalar multiple) by the property $Lf=\alpha f$ and the fact that it is radial. We also observe that $f(0)\neq 0$, because otherwise $f$ would be identically zero.
[…]
We are left with the general case $\alpha\neq\pm 1$. From Observation 4.4 and $Lf=\alpha f$ we get $f(1)=\alpha f(0)$ and $f(2)=\frac{1}{d-1}(d\alpha f(1)-f(0))=\frac{(\alpha^{2}d)-1}{d-1}f(0)$. This implies $f(2)\neq f(0)$: otherwise, we would have $(\alpha^{2}d)-1=d-1$ and thus $\alpha=\pm 1$.
We are ready to prove a major result of this paper, which offers a form of classification of irreducible spherical representations of $G$.
Theorem 4.6.
For any irreducible spherical representation $(\pi,H)$ of $G$ that is not 111-dimensional, there exists $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, such that the representation is equivalent to the left action of $G$ on the subspace of $C$ defined by $\{h\in C:Lh=\alpha h\}$. Conversely, for every $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, there is an irreducible spherical representation of $G$ that results from the construction.
(Note: It is not guaranteed that all these representations preserve a finite-index sesquilinear form; in fact, we will prove later that this is only possible if $\alpha\in\mathbb{R}$)
Proof.
We start by picking a general irreducible spherical representation of $G$ and using the results above. We know that there exist $\alpha\in\mathbb{K}$ and $f\in H\subset C$ such that $Lf=\alpha f$ and the subspace of $K$-invariant functions in $H$ is the space of scalar multiples of $f$. We can assume $\alpha\neq\pm 1$, and thus $f(0)\neq f(2)$: otherwise, as we have seen, the representation is 111-dimensional.
Since the representation is irreducible, $H$ is generated (as a vector space) by $\{\pi(g)f:g\in G\}$. A function $\pi(g)f$ is radial with respect to $go$ and satisfies $L\pi(g)f=\alpha\pi(g)f$. The latter property is preserved under linear combinations, so we have $H\subseteq\{h\in C:Lh=\alpha h\}$.
To obtain the other inclusion, we pick a generic $h\in C$ such that $Lh=\alpha h$ and we show that it is in $H$. Since $H$ is $G$-invariant and $f\in H$, it is enough to show that we can express $h$ as a finite linear combination of translates (by elements of $G$) of $f$.
For every $v\in V$, let $N(v)$ be the set of neighbors of $v$ which are further away from the origin than $v$ (if $v=o$ we take the full set of neighbors). To build $h$ as a finite linear combination of translates of $f$, we start with the function $F=h(o)f(0)^{-1}f$, that agrees with $h$ at the origin. As $LF(o)=\alpha F(o)=\alpha h(o)=Lh(o)$, the sum of $F$ over $N(o)$ agrees with $h$. To make the two functions agree at every vertex in $N(o)$, we add to $F$ the function
$\sum_{w\in N(o)}{(h(w)-F(w))(f(0)-f(2))^{-1}\pi(g_{w})f},$
where $g_{w}\in G$ sends $o$ to $w$. The sum of the coefficients is zero, and this guarantees that $F$ is unchanged at $o$. At every $w\in N(o)$ the function changes by
$(h(w)-F(w))(f(0)-f(2))^{-1}f(0)$
due to the contribution from $\pi(g_{w})f$, and by
$\sum_{z\in N(o)\setminus\{w\}}{(h(z)-F(z))(f(0)-f(2))^{-1}f(2)}=-(h(w)-F(w))(f(0)-f(2))^{-1}f(2)$
due to the contribution from $\pi(g_{z})f$ for $z\neq w$. This adds up to $h(w)-F(w)$, so we have a new function $F$ that agrees with $H$ on the ball $B_{1}(o)$.
We can move forward by applying the same procedure to each of the sets $N(w)$ with $w\in N(o)$: this will allow to update the function $F$ so that it agrees with $h$ inside $N(w)$ without changing the values inside the ball $B_{1}(o)$, and finally get a function that agrees with $h$ on the ball $B_{2}(o)$. We can continue inductively with $B_{3}(o)$ and so on. After finitely many steps $F$ and $h$ will agree everywhere: $h$ is left-invariant under some open compact subgroup that contains $G(B_{n}(o))$ for some $n$ (Proposition 2.2(a)), so it is uniquely determined by the property $Lh=\alpha h$ and the values inside $B_{n}(o)$, which means that it suffices to have $F$ and $h$ agree on $B_{n}(o)$.
We have proved that the space $H$ coincides with $\{h\in C:Lh=\alpha h\}$, so the representation we started from is equivalent to the left action of $G$ on that space.
For the converse, we wish to prove that if we start from a generic $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, and consider the space $\{h\in C:Lh=\alpha h\}$ with the left action of $G$, we have a spherical irreducible representation. We know that there is a radial (hence $K$-invariant) nonzero function in this space, so we are only left to prove irreducibility. Suppose there is a nonzero $G$-invariant subspace $X\subset H$ and pick $h\in X$, $h\neq 0$. It suffices to prove that any function in $H$ can be obtained as a finite linear combination of $G$-translates of $h$.
We have seen before that every function in $H$ is a finite linear combination of $G$-translates of $f$, therefore it suffices to show that $f$ is a finite linear combination of $G$-translates of $h$. Up to translating, we may assume $h(o)\neq 0$. If we average $h$ over $K$ (with respect to the Haar measure), we get a radial function that has to be a nonzero multiple of $f$. The operation of averaging over $K$ is the same as applying a finite linear combination of $G$-translates (actually $K$-translates), since the function $h$ is left-invariant under $G(B_{n}(o))$ for some $n$ and the quotient $K/G(B_{n}(o))=G(o)/G(B_{n}(o))$ is finite.
5 Preserved sesquilinear forms
In this section, we finish the proof of Theorem 2.3 in the case where $G$ acts transitively on $V$. In order to do this, we study nondegenerate sesquilinear forms of finite index that can be preserved by the spherical representations that we have classified in Theorem 4.6.
We first recall that, due to Proposition 3.3, super-cuspidal and special representations (which are unitary) cannot preserve any nondegenerate sesquilinear form of index 1$\geq$ 1≥ 1. This means that to prove Theorem 2.3, it suffices to consider the case of spherical representations.
Therefore, let $(\pi,H)$ be an irreducible spherical representation of $G$ that is not 111-dimensional. We apply Theorem 4.6 and find that there exists $\alpha\in\mathbb{K}$, $\alpha\neq\pm 1$, such that $H$ is the space of functions $h:V\rightarrow\mathbb{K}$ that are left-invariant under some open compact subgroup of $G$ and satisfy $Lh=\alpha h$, and $\pi$ is the natural left action of $G$ on this space.
We split the space $H$ into a direct sum of $K$-invariant subspaces. First, we define $H_{0}$ as the 111-dimensional space generated by the unique (up to scalar multiple) $K$-invariant $f\in H$. For $v\in V$, we define the set $N(v)$ as in the proof of Theorem 4.6: $N(v)$ is the set of neighbors of $v$ that are further away from the origin than $v$ (or the complete set of neighbors if $v=o$). We define the space $\widetilde{H}_{1}$ as the subspace of $H$ consisting of functions that are invariant under $G(B_{1}(o))$, and then $H_{1}$ such that $\widetilde{H}_{1}=H_{0}\oplus H_{1}$, in this way:
$H_{1}:=\left\{h\in\widetilde{H}_{1}:\sum_{v\in N(o)}{h(v)}=0\right\}.$
We can proceed inductively. We define, for $n\in\mathbb{N}$, the set $S_{n}:=\{v\in V:d(o,v)=n\}$. Then we define, for $n\geq 2$, the space $\widetilde{H}_{n}$ of functions in $H$ that are invariant under $G(B_{n}(o))$, and then
$H_{n}:=\left\{h\in\widetilde{H}_{n}:\sum_{w\in N(v)}{h(w)}=0\ \ \ \forall v\in S_{n-1}\right\},$
so that $\widetilde{H}_{n}=\bigoplus_{i=0}^{n}{H_{i}}$.
The space $H$ is the increasing union of $\widetilde{H}_{n}$, and hence is the direct sum of the spaces $H_{n}$. All these spaces are clearly $K$-invariant. We have $\dim(H_{0})=1$, $\dim(H_{1})=d-1$, and $\dim(H_{n})=d(d-2)(d-1)^{n-2}$ for all $n\geq 2$.
It is clear that $H_{0}$ is $K$-irreducible, since it is 111-dimensional. We show that $H_{1}$ is also $K$-irreducible. Since $K$ acts transitively on $\partial T$, it also acts transitively on $S_{1}$. Moreover, if $v\in S_{1}$ the stabilizer of $v$ in $G$ acts transitively on $\partial T$. This implies that its intersection with $K=G(o)$ acts transitively on $S_{1}\setminus\{v\}$, which (since $v$ is generic) is the same as saying that the action of $K$ on $S_{1}$ is doubly transitive. Now, let $X$ be any nonzero $K$-invariant subspace of $H_{1}$ and let $h\in X$ be nonzero. By averaging $h$ over $K\cap G(v)$, where $v\in S_{1}$ satisfies $h(v)\neq 0$, we get a nonzero scalar multiple of a function with value 111 at $v$ and $-\frac{1}{d-1}$ at all other vertices in $S_{1}$. Since $K$-translates of this function span $H_{1}$, we must have $X=H_{1}$.
For spaces $H_{n}$ with $n\geq 2$ we cannot prove $K$-irreducibility, but we can prove the following:
Lemma 5.1.
For every $n\geq 2$, each $K$-irreducible subspace of $H_{n}$ has dimension at least $|S_{n-1}|=d(d-1)^{n-2}$.
Proof.
For $v\in V$, $v\neq o$, we define the cone $C_{v}$ as the set of vertices $w$ such that the unique path between $o$ and $w$ passes through $v$. We define the anti-cone $D_{v}:=(V\setminus C_{v})\cup\{v\}$. As $G$ has the independence property, it follows from Proposition 11 of [1] (applied to the subtree $B_{n-1}(o)$) that
$G(B_{n-1}(o))=\prod_{v\in S_{n-1}}G(D_{v}).$
Clearly, all factors commute. If $X$ is a $K$-irreducible subspace of $H_{n}$ and $h\in X$ is nonzero, we can choose $v\in S_{n-1}$ such that $h(w)\neq 0$ for at least one $w\in N(v)$. We average $h$ across $\prod_{z\in S_{n-1}\setminus\{v\}}G(D_{v})$ and get a function that is unchanged on $N(v)$ and zero on $N(z)$ for every $z\in S_{n-1}\setminus\{v\}$. By the transitivity of $K$ on $S_{n-1}$ we can then get such a function for any $v\in S_{n-1}$. In this way, we get $|S_{n-1}|$ linearly independent functions in $X$, which proves our claim.
Since $|S_{n-1}|=d(d-1)^{n-2}>\dim{H_{n-1}}$ for every $n\geq 2$, the lemma implies the following:
Corollary 5.2.
Let $i,j\in\mathbb{N}$ with $i\neq j$. Then, no $K$-irreducible subspace of $H_{i}$ can be isomorphic to any $K$-irreducible subspace of $H_{j}$.
To continue the proof of Theorem 2.3, we fix a $G$-invariant sesquilinear form $B$ on $H$. It follows from the Schur Lemma and Corollary 5.2 that $B$ must be the sum (for $n\in\mathbb{N}$) of $K$-invariant sesquilinear forms $B_{n}$ on $H_{n}$. This implies that if $\mathcal{A}$ and $\mathcal{B}$ are disjoint subsets of $\mathbb{N}$ then every function in $\bigoplus_{n\in\mathcal{A}}{H_{n}}$ has to be orthogonal to every function in $\bigoplus_{n\in\mathcal{B}}{H_{n}}$. We can now show the following:
Lemma 5.3.
If the representation preserves the finite-index sesquilinear form $B$ (which we assume to be antilinear in the first variable and linear in the second), then $\alpha\in\mathbb{R}$.
Proof.
Let $f$ be the unique function in $H_{0}$ satisfying $f(0)=1$. It satisfies $f(1)=\alpha$ and $f(2)=\frac{(\alpha^{2}d)-1}{d-1}$. Let $h$ be a function in $H_{1}$ satisfying $h(v)=1$ for one particular $v\in N(o)$ and $h(w)=-\frac{1}{d-1}$ for all $w\in N(o)\setminus\{v\}$. Since $f\in H_{0}$ and $h\in H_{1}$, we have $B(f,h)=0$.
If we consider $g\in G$ that sends $o$ to $v$ and $v$ to $o$, we have $\pi(g)f=\alpha f+(1-\alpha^{2})h$ and $\pi(g)h=f-\alpha h$. The latter implies
$B(f,f)=\left(1-|\alpha|^{2}\right)B(h,h).$
Observe that $B(f,f)\neq 0$, because otherwise the form would be degenerate ($f$ is orthogonal to all functions in $\bigoplus_{n\geq 1}{H_{n}}$). Hence the last relation implies $B(h,h)\neq 0$.
As $B(f,h)=0$ it must be $B(\pi(g)f,\pi(g)h)=0$. This means that $B(\alpha f+(1-\alpha^{2})h,f-\alpha h)=0$, and therefore
$\overline{\alpha}B(f,f)-(1-\overline{\alpha}^{2})\alpha B(h,h)=0.$
Combining this with the expression we got before for $B(f,f)$ we get
$\overline{\alpha}\left(1-|\alpha|^{2}\right)B(h,h)=(1-\overline{\alpha}^{2})\alpha B(h,h).$
We can divide by $B(h,h)$ on the right as it is nonzero. By developing the two expressions we get
$\overline{\alpha}-\overline{\alpha}|\alpha|^{2}=\alpha-\overline{\alpha}|\alpha|^{2}$
and therefore $\overline{\alpha}=\alpha$, meaning that $\alpha\in\mathbb{R}$.
Corollary 5.4.
Every spherical irreducible complex or quaternionic representation of $G$ preserving a finite-index nondegenerate sesquilinear form is obtained by tensoring with $\mathbb{C}$ or $\mathbb{H}$ a spherical irreducible real representation of $G$ preserving a finite-index nondegenerate bilinear form.
Therefore, from now on, we may assume $\mathbb{K}=\mathbb{R}$, and we are left to show that no nondegenerate bilinear form $B$ of finite index >1>1> 1 can be preserved by a spherical representation of $G$.
The strategy will be to present, for every $\alpha\in\mathbb{R}\setminus\{\pm 1\}$, an explicit bilinear form of index 0 or 111 that is preserved by the corresponding representation of $G$. This will exclude, by Proposition 3.3, that the representation can preserve a bilinear form of index >1>1> 1, proving the theorem. Furthermore, this will provide a classification of bilinear forms that can be preserved by a spherical irreducible representation. As we have said, from now on we work with real representations, but the same results will hold for complex and quaternionic representations.
Recall that for every $n\geq 1$, a function in $H_{n}$ is uniquely determined by its values on the set of vertices $S_{n}=\{v\in V:d(o,v)=n\}$. More precisely, each function $S_{n}\rightarrow\mathbb{R}$ with sum 0 on the set of neighbors of any vertex in $S_{n-1}$ can be extended in a unique way to a function in $H_{n}$.
We observe that if we define the bilinear form $Q_{n}$ on $H_{n}$ ($n\geq 1$) so that it coincides with the restriction of the standard scalar product on the space of functions $S_{n}\rightarrow\mathbb{R}$ to the subspace consisting of functions with sum 0 on the set of neighbors of any vertex in $S_{n-1}$, then it is preserved by the action of $K=G(o)$.
Let $Q_{0}$ be the bilinear form on $H_{0}$ defined by $Q_{0}(f,f)=1$ (where $f\in H_{0}$ satisfies $f(0)=1$). We define the bilinear form $Q$ on $H$ as follows:
$Q:=\frac{d}{d-1}(1-\alpha^{2})Q_{0}+\sum_{n\geq 1}{Q_{n}}.$
This bilinear form has index 0 if $|\alpha|<1$ and index 111 if $|\alpha|>1$.
To finish the proof of Theorem 2.3 in the case where $G$ acts transitively on $V$, it suffices to show the following:
Proposition 5.5.
The bilinear form $Q$ is preserved under $G$.
It is clear from the definition that $Q$ is preserved under $K$.
Fix a neighbor $v$ of $o$ and consider $g\in G$ that sends $o$ to $v$ and $v$ to $o$.
Lemma 5.6.
The element $g$ and the subgroup $K$ generate the group $G$.
Proof.
Let $g_{0}\in G$. To show that $g_{0}$ can be generated with $g$ and $K$, let $w:=g_{0}^{-1}(o)$. If we find $g_{1}$ generated with $g$ and $K$ such that $g_{1}^{-1}(o)=w$, we will have $g_{0}g_{1}^{-1}(o)=o$ and therefore $g_{0}g_{1}^{-1}\in K$, concluding.
To find such a $g_{1}$, let $m:=d(o,w)$. Observe that by composing $g$ with some element of $K$ we can get an element $g^{\prime}$ that sends $o$ to $v$ and $v$ to a vertex in $S_{2}$. Then the element $(g^{\prime})^{m}$ will send $o$ to a vertex in $S_{m}$, and by composing with an element of $K$ we can get $g^{\prime\prime}$ such that $g^{\prime\prime}(o)=w$. Now, it suffices to take $g_{1}=(g^{\prime\prime})^{-1}$.
The lemma implies that, to prove Proposition 5.5, it suffices to show that $Q$ is preserved under $g$.
Define the functions $f\in H_{0}$ and $h\in H_{1}$ as in the proof of Lemma 5.3. From the definition of $Q$ we have $Q(f,f)=\frac{d}{d-1}(1-\alpha^{2})$ and $Q(h,h)=\frac{d}{d-1}$, since the $d$-dimensional vector $(1,-\frac{1}{d-1},\cdots,-\frac{1}{d-1})$ has Euclidean norm $\frac{d}{d-1}$. Hence $Q(f,f)=(1-\alpha^{2})Q(h,h)$, and therefore $g$ preserves the norm of $h$ and the orthogonality of $f$ and $h$: the two conditions were explicitated in the proof of Lemma 5.3, and it is immediate to see that both hold if $\alpha\in\mathbb{R}$ and $B(f,f)=(1-\alpha^{2})B(h,h)$. The norm of $f$ is also preserved, because $\pi(g)f=\alpha f+(1-\alpha^{2})h$ and the norm of this function is
$\alpha^{2}Q(f,f)+(1-\alpha^{2})^{2}Q(h,h)=(\alpha^{2}+(1-\alpha^{2}))Q(f,f)=Q(f,f).$
We have shown that $g$ preserves $Q$ in $\mathrm{Span}\{f,h\}$. Observe now that the orthogonal complement of $\mathrm{Span}\{f,h\}$ in $H$ is the space of functions in $H$ with value 0 at $o$ and $v$. In fact, it is clear that all such functions are orthogonal to $f$ and $h$, and to see that they generate $H$ together with $f$ and $h$, it suffices to take any $t\in H$ and subtract $t(o)f+(t(v)-\alpha t(o))h$ to obtain a function that vanishes at $o$ and $v$. Furthermore, since $g$ swaps $o$ and $v$, the space $\mathrm{Span}\{f,h\}^{\perp}$ is invariant under $g$.
Therefore, to complete the proof of Proposition 5.5, it suffices to show that $g$ preserves the norm of every function in $\mathrm{Span}\{f,h\}^{\perp}$. To see this, we further split each $H_{n}$ ($n\geq 2$) into two orthogonal subspaces: $H_{n}=H_{n}^{o}\oplus H_{n}^{v}$, where $H_{n}^{o}$ is the space of functions in $H_{n}$ supported on $\{w\in V:d(o,w)<d(v,w)\}$, and $H_{n}^{v}$ is the space of functions in $H_{n}$ supported on $\{w\in V:d(o,w)>d(v,w)\}$. We can also split $H_{1}$ as $H_{1}=H_{1}^{o}\oplus\mathrm{Span}\{h\}$. For every $j\in\mathrm{Span}\{f,h\}^{\perp}$, we call $j_{n}^{o}$ its projection to $H_{n}^{o}$ and $j_{n}^{v}$ its projection to $H_{n}^{v}$. We have
$\pi(g)j=\pi(g)\left(\sum_{n\geq 1}{j_{n}^{o}}+\sum_{n\geq 2}{j_{n}^{v}}\right)=\sum_{n\geq 1}{\pi(g)j_{n}^{o}}+\sum_{n\geq 2}{\pi(g)j_{n}^{v}}.$
For every $n\geq 1$ we have $\pi(g)j_{n}^{o}\in H_{n+1}^{v}$, and for every $n\geq 2$ we have $\pi(g)j_{n}^{v}\in H_{n-1}^{o}$. From the definition of the bilinear form $Q$ it is immediately seen that the norm of each $j_{n}^{o}$ and $j_{n}^{v}$ is preserved, and since all addends are orthogonal it follows that the norm of $j$ is preserved.
This completes the proof of Theorem 2.3 in the case where $G$ acts transitively on $V$.
6 The non-transitive case
In this section, we generalize the results of the previous two sections to the case where the action of $G$ on $V$ is not transitive. In particular, this will lead to a complete proof of Theorem 2.3.
We recall that, due to Proposition 2.2(c), if the action of $G$ on $V$ is not transitive, then it has two orbits, one consisting of vertices $v$ with $d(o,v)$ even and one of those with $d(o,v)$ odd.
The tree must be homogeneous or semi-homogeneous. We may assume that it is semi-homogeneous of degrees $r,s\geq 3$: the homogeneous case will be the special case $r=s$. We may also assume, without loss of generality, that the base vertex $o$ has degree $r$.
We proceed as in Sections 4 and 5, adapting the results to this new case.
Proposition 3.2 implies that $(\pi,H)$ is equivalent to a subrepresentation of $(\Pi,C)$, where $C$ is the space of functions on $W:=\{v\in V:d(o,v)\text{ even}\}$ that are left-invariant under some open compact subgroup of $G$, and $\Pi$ is the left action of $G$ on this space.
We cannot use the standard Laplacian defined in Definition 4.3, so we introduce:
Definition 6.1.
(2-Laplacian) The 2-Laplacian on the tree $T$ is the operator $L_{2}:C\rightarrow C$ defined as follows: if $f:W\rightarrow\mathbb{K}$ is a function in $C$, the function $L_{2}f$ associates to every vertex $v$ in $W$ the average of the values of $f$ at vertices $w$ satisfying $d(v,w)=2$.
Observation 4.4 changes in the following way: the 2-Laplacian of a radial function is again a radial function, with values
  • $L_{2}f(0)=f(2)$,
  • $L_{2}f(2n)=\frac{1}{r(s-1)}f(2n-2)+\frac{s-2}{r(s-1)}f(2n)+\frac{r-1}{r}f(2n+2)$ for all $n\geq 2$.
This implies, similarly to the transitive case with the Laplacian, that for every $\alpha\in\mathbb{K}$ the space of radial functions that satisfy $L_{2}f=\alpha f$ is 111-dimensional.
Analogously to the transitive case, we can use the Gel’fand pair property of $(G,K)$ to show that there exists a function $f\in H$ such that the space of $K$-invariant functions in $H$ is the space of scalar multiples of $f$, and there exists $\alpha\in\mathbb{K}$ such that $L_{2}f=\alpha f$.
We can see that there is one special case, namely $\alpha=1$, where the function $f$ is constant on $W$ and therefore $(\pi,H)$ is the trivial representation. In all other cases, we have $f(0)\neq f(2)$. We wish to deduce an analog of Theorem 4.6. It turns out that the natural analog is always true, except in one particular case.
Theorem 6.2.
If $\alpha\neq 1$ and $\alpha\neq-\frac{1}{s-1}$, then the natural analog of Theorem 4.6 holds: the representation $(\pi,H)$ is equivalent to the left action of $G$ on the space $\{h\in C:L_{2}h=\alpha h\}$, which is spanned by the $G$-translates of $f$, and conversely for every $\alpha\in\mathbb{K}$, $\alpha\neq 1$, $\alpha\neq-\frac{1}{s-1}$, the construction gives an irreducible spherical representation of $G$.
Proof.
We can repeat the proof of Theorem 4.6. The only part that needs to be changed (and that will lead to the exception for $\alpha=-\frac{1}{s-1}$) is when we show that the space $\{h\in C:L_{2}h=\alpha h\}$ is generated by the $G$-translates of the radial function $f$. To show this in our case, we fix a function $h$ in the space and try to build it with scalar multiples of translates of $f$.
We start with $F=h(o)f(0)^{-1}f$, which agrees with $h$ at $o$. Now we wish to make it agree with $h$ on $S_{2}$ as well. For every $w\in S_{2}$, let $g_{w}\in G$ send $o$ to $w$. We add to $F$ the function
$\sum_{v\in N(o)}{\sum_{w\in N(v)}{(h(w)-F(w))(f(0)-f(2))^{-1}\pi(g_{w})f}}.$
The sum of the coefficients is zero, which guarantees that $F$ is unchanged at $o$. For every $v\in N(o)$, let $R(v):=\sum_{w\in N(v)}{(h(w)-F(w))}$. The condition $L_{2}h=\alpha h$, together with $L_{2}F=\alpha F$, implies that $\sum_{v\in N(o)}{R(v)}=0$.
For every $v\in N(o)$ and for every $w\in N(v)$, the function at $w$ has been changed by
$(h(w)-F(w))(f(0)-f(2))^{-1}f(0)$
due to the contribution from $\pi(g_{w})f$, by
$\sum_{z\in N(v)\setminus\{w\}}{(h(w)-F(w))(f(0)-f(2))^{-1}}f(2)=(R(v)-(h(w)-F(w)))(f(0)-f(2))^{-1}f(2)$
due to the contribution from $\pi(g_{z})f$ with $z\in N(v)\setminus\{w\}$, and by
$\sum_{x\in N(o)\setminus\{v\}}{\sum_{y\in N(x)}{(h(y)-F(y))(f(0)-f(2))^{-1}}}f(4)=-R(v)(f(0)-f(2))^{-1}f(4)$
due to the contribution from $\pi(g_{y})f$ with $y\in N(x)$ where $x\in N(o)\setminus\{v\}$.
All this adds up to $h(w)-F(w)+R(v)(f(0)-f(2))^{-1}(f(2)-f(4))$.
To remove the extra term $T(v):=R(v)(f(0)-f(2))^{-1}(f(2)-f(4))$, we subtract from $F$ the function
$\sum_{v\in N(o)}{T(v)(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}}\sum_{w\in N(v)}{\pi(g_{w})f}.$
We need to assume that $f(0)+(s-2)f(2)-(s-1)f(4)\neq 0$. At every $w$ the value will be changed by
$-T(v)(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}f(0)$
due to the contribution from $\pi(g_{w})f$, by
$-T(v)(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}(s-2)f(2)$
due to the contribution from $\pi(g_{z})f$ with $z\in N(v)\setminus\{w\}$, and by
$\sum_{x\in N(o)\setminus\{v\}}{-T(x)(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}(s-1)f(4)}=$
$=T(v)(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}(s-1)f(4)$
due to the contribution from $\pi(g_{y})f$ with $y\in N(x)$ where $x\in N(o)\setminus\{v\}$.
All this adds up to $-T(v)$, so now the function $F$ agrees with $h$ on $S_{2}$ as well as at $o$.
Like in the proof of Theorem 4.6, we can follow the same procedure to make the two functions also agree on $S_{4}$ (applying the procedure to all vertices in $S_{2}$ in place of $o$), then on $S_{6}$ and so on, getting them to agree everywhere in a finite number of steps since $h$ is invariant under an open compact subgroup of $G$.
Therefore, the theorem is proved in the non-transitive case whenever $f(0)+(s-2)f(2)-(s-1)f(4)\neq 0$. As $f(2)=\alpha f(0)$ and
$f(4)$ $=\frac{r}{r-1}\left(\left(\alpha-\frac{s-2}{r(s-1)}\right)f(2)-\frac{1}{r(s-1)}f(0)\right)=$
$=\frac{r(s-1)\alpha^{2}-(s-2)\alpha-1}{(r-1)(s-1)}f(0),$
the condition is true if and only if $1+(s-2)\alpha-\frac{r(s-1)\alpha^{2}-(s-2)\alpha-1}{(r-1)}\neq 0$, which is equivalent to $(s-1)\alpha^{2}-(s-2)\alpha-1\neq 0$, i. e. $((s-1)\alpha+1)(\alpha-1)\neq 0$, which holds for all $\alpha\neq 1,-\frac{1}{s-1}$.
We are left with the case $\alpha=-\frac{1}{s-1}$. In that case, the function $f$ satisfies $f(2n)=\alpha^{n}f(0)$ for all $n\in\mathbb{N}$. This is checked directly for $n=0,1$ and by induction for $n\geq 2$: we have
$f(2n)$ $=\frac{r}{r-1}\left(\left(\alpha-\frac{s-2}{r(s-1)}\right)f(2n-2)-\frac{1}{r(s-1)}f(2n-4)\right)=$
$=\frac{r}{r-1}\left(1+\frac{s-2}{r}-\frac{s-1}{r}\right)\alpha^{n}=\alpha^{n}.$
Theorem 6.3.
If $\alpha=-\frac{1}{s-1}$, the space $H$ is the subspace of $\{h\in C:L_{2}h=\alpha h\}$ consisting of functions with sum zero on the set of neighbors of any vertex in $V\setminus W$ (i. e. any vertex at odd distance from $o$).
Proof.
[…]
We only need to show that the $G$-translates of $f$ span our space. We can follow the exact same proof of Theorem 6.2, up to the point where we are left with the term $T(v)=R(v)(f(0)-f(2))^{-1}(f(2)-f(4))$. We cannot eliminate this term with the procedure that follows in the proof, because it would involve multiplying functions by the inverse of zero. However, we can show that in our case there is nothing to eliminate, as we have $R(v)=0$.
Recall that, in fact, $R(v)$ was the sum of $h(w)-F(w)$ where $w$ varies in the set $N(v)$. If we let $\overline{v}$ be the unique neighbor of $v$ which is not in $N(v)$ (i. e. the one on the path to $o$), as the sum of $h-F$ over $N(v)\cup\{\overline{v}\}$ is zero we have $R(v)=F(\overline{v})-h(\overline{v})$. The functions $F$ and $h$ already agree at $\overline{v}$, therefore $R(v)=0$.
We can proceed to adapting the proof of Theorem 2.3 to the non-transitive case. We first assume $\alpha\neq-\frac{1}{s-1}$; the remaining case will be treated in the end.
We divide $H=\{h\in C:L_{2}h=\alpha h\}$ into a direct sum of $K$-invariant subspaces: $H=\bigoplus_{n\in\mathbb{N}}{H_{n}}$. Again, $H_{0}$ is 111-dimensional (spanned by a radial $f$). Then, for $n\geq 1$ we define $\widetilde{H}_{n}$ as the subspace of $H$ consisting of functions that are invariant under $G(B_{n}(o))$. In this way, a function in $\widetilde{H}_{n}$ will be determined by its values in $S_{2k}$ if $n\in\{2k-1,2k\}$, and in the case $n=2k-1$ the value will have to be constant on $N(w)$ for every $w\in S_{2k-1}$. We can define, for every $k\geq 1$:
$H_{2k-1}:=\left\{h\in\widetilde{H}_{2k-1}:\sum_{w\in N(v)}{\sum_{x\in N(w)}{h(x)}}=0\ \forall v\in S_{2k-2}\right\};$
$H_{2k}:=\left\{h\in\widetilde{H}_{2k}:\sum_{x\in N(w)}{h(x)}=0\ \forall w\in S_{2k-1}\right\}.$
In this way, we have $\widetilde{H}_{n}=\bigoplus_{i=0}^{n}{H_{i}}$ for every $n$, and $H=\bigoplus_{n\in\mathbb{N}}{H_{n}}$. The dimensions of the subspaces $H_{n}$ are as follows:
  • $\dim(H_{0})=1$,
  • $\dim(H_{1})=r-1$,
  • $\dim(H_{2})=r(s-2)$,
  • $\dim(H_{2k-1})=r(r-2)(r-1)^{k-2}(s-1)^{k-1}$ for $k\geq 2$,
  • $\dim(H_{2k})=r(r-1)^{k-1}(s-2)(s-1)^{k-1}$ for $k\geq 2$.
Like in the transitive case, we prove that $H_{1}$ is irreducible and that for every $n\geq 2$ each $K$-irreducible subspace of $H_{n}$ has dimension at least $|S_{n-1}|$. The proof is exactly the same if $n$ is even, while if $n=2k-1$ is odd, we consider, in place of functions on $S_{2k-1}$, functions on $S_{2k}$ that are constant on $N(w)$ for every $w\in S_{2k-1}$ (as is the case for functions in $H_{2k-1}$).
Since for every $k\geq 1$ we have $|S_{2k-1}|=r(r-1)^{k-1}(s-1)^{k-1}$ and $|S_{2k}|=r(r-1)^{k-1}(s-1)^{k}$, we always have $|S_{n-1}|>\dim{H_{n-1}}$. Hence, Corollary 5.2 holds, and therefore if the representation preserves a sesquilinear form $B$ then it must be the sum (for $n\in\mathbb{N}$) of $K$-invariant sesquilinear forms $B_{n}$ on $H_{n}$.
Lemma 6.4.
If the representation preserves the finite-index sesquilinear form $B$ (which we assume to be antilinear in the first variable and linear in the second), then $\alpha\in\mathbb{R}$.
Proof.
Let $f$ be the unique function in $H_{0}$ that satisfies $f(0)=1$. It satisfies $f(2)=\alpha$ and $f(4)=\frac{r(s-1)\alpha^{2}-(s-2)\alpha-1}{(r-1)(s-1)}$. Let $h$ be a function in $H_{1}\oplus H_{2}$ such that there exist $v\in N(o)$ and $x\in N(v)$ with $h(x)=1+\alpha$, $h(w)=\alpha$ for all $w\in N(v)\setminus\{x\}$, and $h(z)=-\frac{1+(s-1)\alpha}{(r-1)(s-1)}$ for all $z\in S_{2}\setminus N(v)$.
If we consider $g\in G$ that sends $o$ to $x$ and $x$ to $o$, we have $\pi(g)f=\alpha f+(1-\alpha)h$ and $\pi(g)h=(1+\alpha)f-\alpha h$. The former implies
$|1-\alpha|^{2}B(h,h)=\left(1-|\alpha|^{2}\right)B(f,f).$
Observe that $B(f,f)\neq 0$, because otherwise the form would be degenerate ($f$ is orthogonal to all functions in $\bigoplus_{n\geq 1}{H_{n}}$).
Since $B(f,h)=0$, it must be $B(\pi(g)f,\pi(g)h)=0$. This means that $B(\alpha f+(1-\alpha)h,(1+\alpha)f-\alpha h)=0$, and therefore
$\overline{\alpha}(1+\alpha)B(f,f)-(1-\overline{\alpha})\alpha B(h,h)=0.$
Multiplying by $1-\alpha$ on the left and using the equality above, we get
$(1-\alpha)\overline{\alpha}(1+\alpha)B(f,f)=|1-\alpha|^{2}\alpha B(h,h)=\alpha\left(1-|\alpha|^{2}\right)B(f,f).$
We can divide by $B(f,f)$ on the right, since it is nonzero. By developing the two expressions, we get
$\overline{\alpha}-\alpha|\alpha|^{2}=\alpha-\alpha|\alpha|^{2}$
and therefore $\overline{\alpha}=\alpha$, which means that $\alpha\in\mathbb{R}$.
Lemma 6.4 implies that Corollary 5.4 also holds in the non-transitive case.
Like in the transitive case, we finish the proof of Theorem 2.3 by presenting, for every $\alpha\in\mathbb{R}\setminus\{1\}$, an explicit bilinear form preserved under $G$. In the non-transitive case, the bilinear form will be of index 0, 111 or $\infty$: in the first two cases, Proposition 3.3 applies directly and excludes that $G$ can preserve a form of index >1>1> 1, while in the third case we can argue by contradiction that if $G$ preserved a form of finite index $p$ it could not preserve (again by Proposition 3.3) a form of index $\infty$.
For every $n\in\mathbb{N}$, a function in $H_{n}$ is uniquely determined by its values in $S_{\lceil\frac{n}{2}\rceil}$. We define the bilinear form $Q_{n}$ on $H_{n}$ in such a way that it coincides with the restriction of the standard scalar product on the space of functions $S_{\lceil\frac{n}{2}\rceil}\rightarrow\mathbb{R}$ to the subspace consisting of those that give rise to a function in $H_{n}$.
We define the bilinear form $Q$ on $H$ as follows:
$Q:=(1-\alpha)Q_{0}+\frac{r-1}{r(1+(s-1)\alpha)}\sum_{k\geq 1}{Q_{2k-1}}+\sum_{k\geq 1}{Q_{2k}}.$
This bilinear form has index 0 if $-\frac{1}{s-1}<\alpha<1$, index 111 if $\alpha>1$, and index $\infty$ if $\alpha<-\frac{1}{s-1}$.
We need to check that this bilinear form is preserved under $G$. Let $v\in N(o)$ and let $w\in N(v)$. Consider $g\in G$ that sends $o$ to $w$ and $w$ to $o$. The proof of Lemma 5.6 can be easily adapted to show that $g$ and $K$ generate $G$. As the bilinear form $Q$ is clearly preserved under $K$, it suffices to show that it is preserved under $g$.
Let $D:=\{o\}\cup N(v)$ be the set of neighbors of $v$. The space $H$ can be split into two subspaces: the space $H_{D}$ of functions that are invariant under $G(D)$, and the space $H_{D}^{\perp}$ of functions that are zero on $D$. We have $H=H_{D}\oplus H_{D}^{\perp}$: the intersection is easily seen to be trivial, and for every $h\in H$ there is a $G(D)$-invariant function that agrees with it on $D$ (built with translates of the radial function $f$ by elements of $G$ that fix $v$) and whose difference with $h$ will thus be zero on $D$. The notation $H_{D}^{\perp}$ is not casual, as it is easy to see that every function in $H_{D}$ is orthogonal to every function in $H_{D}^{\perp}$ (with respect to the bilinear form $Q$). It is also easy to see that the two subspaces are preserved under $g$ ($g$ preserves $D$), therefore it suffices to show that $g$ preserves $Q$ on $H_{D}$ and on $H_{D}^{\perp}$.
We first consider $H_{D}$. We split it into the direct sum $H_{D}^{(1)}\oplus H_{D}^{(2)}\oplus H_{D}^{(3)}$, where:
$H_{D}^{(1)}:=\left\{h\in H_{D}:h(o)=h(w),\ h\text{ constant on }D\setminus\{o,w\}\right\};$
$H_{D}^{(2)}:=\left\{j\in H_{D}:j(o)+j(w)=0,\ j(x)=0\ \forall x\in D\setminus\{o,w\}\right\};$
$H_{D}^{(3)}=\left\{l\in H_{D}:l(0)=l(w)=0,\ \sum_{x\in D\setminus\{o,w\}}{l(x)}=0\right\}.$
The three subspaces are preserved by $g$, as $g$ swaps $o$ and $w$.
Let $h,j,l$ be arbitrary functions in $H_{D}^{(1)},H_{D}^{(2)},H_{D}^{(3)}$ respectively. We have $\pi(g)h=h$, therefore $g$ preserves $Q$ on $H_{D}^{(1)}$. We also have $\pi(g)j=-j$, therefore $g$ preserves $Q$ on $H_{D}^{(2)}$. The functions $l$ and $\pi(g)l$ are both in $H_{2}$ with the same norm, therefore $g$ preserves $Q$ on $H_{D}^{(3)}$.
We claim that the spaces $H_{D}^{(1)}$, $H_{D}^{(2)}$, $H_{D}^{(3)}$ are mutually orthogonal with respect to the bilinear form $Q$, which will allow us to conclude that $g$ preserves $Q$ on $H_{D}$. For this, we need to show that $h,j,l$ are mutually orthogonal.
First, we see that $l$ is orthogonal to both $h$ and $j$, because its restriction to $S_{2}$ is supported (with sum zero) on $N(v)\setminus\{w\}$ while $h$ and $j$ are constant on this set.
Now, we need to show that $Q(h,j)=0$. We may assume (up to multiplying by a scalar) that $j(o)=1$, which implies $j(w)=-1$. Let $a$ be the value of $h$ at $o$ (or $w$) and let $b$ be its value at any $x\in D\setminus\{o,w\}$. The function $h$ is the sum of the following:
  • a function in $H_{0}$ with value $a$ at $o$,
  • a function in $H_{1}$ with value $\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b$ at any $x\in N(v)$ and value $-\frac{1}{r-1}\left(\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b\right)$ at any $x\in S_{2}\setminus N(v)$,
  • a function in $H_{2}$ supported on $N(v)$ with value $\frac{s-2}{s-1}(a-b)$ at $w$ and value $-\frac{1}{s-1}(a-b)$ at any $x\in N(v)\setminus\{w\}$.
The function $j$ is the sum of the following:
  • a function in $H_{0}$ with value 111 at $o$,
  • a function in $H_{1}$ with value $-\frac{1}{s-1}-\alpha$ at any $x\in N(v)$ and value $\frac{1}{r-1}\left(\frac{1}{s-1}+\alpha\right)$ at any $x\in S_{2}\setminus N(v)$,
  • a function in $H_{2}$ supported on $N(v)$ with value $-\frac{s-2}{s-1}$ at $w$ and $\frac{1}{s-1}$ at any $x\in N(v)\setminus\{w\}$.
We have:
$Q(h,j)$ $=(1-\alpha)a+\frac{r-1}{r(1+(s-1)\alpha)}(s-1)\left(\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b\right)\left(-\frac{1}{s-1}-\alpha\right)+$
$+\frac{r-1}{r(1+(s-1)\alpha)}(r-1)(s-1)\left(-\frac{1}{r-1}\left(\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b\right)\right)\frac{1}{r-1}\left(\frac{1}{s-1}+\alpha\right)+$
$+\frac{s-2}{s-1}(a-b)\left(-\frac{s-2}{s-1}\right)+(s-2)\left(-\frac{1}{s-1}(a-b)\right)\frac{1}{s-1}=$
$=(1-\alpha)a-\frac{r-1}{r(1+(s-1)\alpha)}\left(\left(1-(s-1)\alpha\right)a+(s-2)b\right)\left(\frac{1}{s-1}+\alpha\right)+$
$-\frac{1}{r(1+(s-1)\alpha)}\left(\left(1-(s-1)\alpha\right)a+(s-2)b\right)\left(\frac{1}{s-1}+\alpha\right)-\frac{(s-2)^{2}+(s-2)}{(s-1)^{2}}(a-b)=$
$=(1-\alpha)a-\frac{1}{s-1}\left(\left(1-(s-1)\alpha\right)a+(s-2)b\right)-\frac{s-2}{s-1}(a-b)=$
$=\frac{1}{s-1}(((1-\alpha)(s-1)-(1-(s-1)\alpha)-(s-2))a-((s-2)-(s-2))b)=0.$
Therefore, we have shown that $g$ preserves $Q$ on $H_{D}$. We now show that $g$ preserves $Q$ on $H_{D}^{+}$, in a way similar to the end of the proof of Proposition 5.5.
We split each $H_{n}$ ($n\geq 3$) in three orthogonal subspaces: $H_{n}=H_{n}^{o}\oplus H_{n}^{x}\oplus H_{n}^{w}$, where $H_{n}^{o}$ is the space of functions in $H_{n}$ supported on $\{z\in W:d(o,z)<d(w,z)\}$ (which means that the path from $z$ to $v$ meets $D$ at $o$), $H_{n}^{x}$ is the space of functions in $H_{n}$ supported on $\{z\in W:d(o,z)=d(w,z)\}$ (which means that the path from $z$ to $v$ meets $D$ at a vertex other than $o$ and $w$), and $H_{n}^{w}$ is the space of functions in $H_{n}$ supported on $\{z\in W:d(o,z)>d(w,z)\}$ (which means that the path from $z$ to $v$ meets $D$ at $w$). We can also define $H_{1}^{o}$ as the space of functions in $H_{1}$ supported on $\{z\in W:d(o,z)<d(v,z)\}$, and $H_{2}^{o}$ in the same way. Observe that a function in $H_{D}^{\perp}$ lies in $H_{1}^{o}\oplus H_{2}^{o}\oplus\bigoplus_{n\geq 3}{H_{n}}$.
For every $q\in H_{D}^{\perp}$, we call $q_{n}^{o}$ its projection to $H_{n}^{o}$, $q_{n}^{x}$ its projection to $H_{n}^{x}$, and $q_{n}^{w}$ its projection to $H_{n}^{w}$. We have
$\pi(g)q=\pi(g)\left(\sum_{n\geq 1}{q_{n}^{o}}+\sum_{n\geq 3}{q_{n}^{x}}+\sum_{n\geq 3}{q_{n}^{w}}\right)=\sum_{n\geq 1}{\pi(g)q_{n}^{o}}+\sum_{n\geq 3}{\pi(g)q_{n}^{x}}+\sum_{n\geq 3}{\pi(g)q_{n}^{w}}.$
For every $n\geq 1$ we have $\pi(g)q_{n}^{o}\in H_{n+2}^{w}$, for every $n\geq 3$ we have $\pi(g)q_{n}^{x}\in H_{n}^{x}$, and for every $n\geq 3$ we have $\pi(g)q_{n}^{w}\in H_{n-2}^{o}$. From the definition of the bilinear form $Q$ it is then immediate to see that the norm of each $q_{n}^{o}$, $q_{n}^{x}$ and $q_{n}^{w}$ is preserved, and as all addends are orthogonal, it follows that the norm of $q$ is preserved. This completes the proof of Theorem 2.3 in the non-transitive case, if we assume $\alpha\neq-\frac{1}{s-1}$.
Therefore, we are only left with the case $\alpha=-\frac{1}{s-1}$.
We claim that, in this case, all spaces $H_{2k-1}$ ($k\geq 1$) are trivial. Recall that if $\alpha=-\frac{1}{s-1}$, every function in $H$ has sum zero on the set of neighbors of any vertex at odd distance from $o$. Recall also that a function in $H_{2k-1}$ is zero on $B_{2k-2}(o)$ and constant on $N(v)$ for every $v\in S_{2k-1}$. As the sum over neighbors of $v$ has to be zero, the sum over $N(v)$ has to be zero, which means that the function is identically zero on $N(v)$. This shows that the function is identically zero on $S_{2k}$. Then, since the function is invariant under $G(B_{2k-1})$, it has to be zero everywhere.
For every even $n$ we can define the bilinear form $Q_{n}$ on $H_{n}$ as in the case $\alpha\neq-\frac{1}{s-1}$, then we can define the bilinear form $Q$ on $H$ as follows:
$Q:=(1-\alpha)Q_{0}+\sum_{k\geq 1}{Q_{2k}}.$
This form has index 0 (it is positive definite). To verify that it is preserved under $G$, the proof is the same as in the case $\alpha\neq-\frac{1}{s-1}$, only the verification that $g$ preserves $Q$ on $H_{D}$ must be adapted.
[…]
The proof that $g$ preserves the three subspaces, that it preserves $Q$ on each of them, and that $H_{D}^{(3)}$ is orthogonal to both of the others works exactly as in the case $\alpha\neq-\frac{1}{s-1}$. The only point to change is where we check that $H_{D}^{(1)}$ and $H_{D}^{(2)}$ are orthogonal.
[…]
$Q(h,j)$ $=(1-\alpha)a+\frac{s-2}{s-1}(a-b)\left(-\frac{s-2}{s-1}\right)+(s-2)\left(-\frac{1}{s-1}(a-b)\right)\frac{1}{s-1}=$
$=\left(1+\frac{1}{s-1}\right)a-\frac{s-2}{s-1}(a-b)=\frac{1}{s-1}(2a+(s-2)b)=0.$
This concludes the proof for the case $\alpha=-\frac{1}{s-1}$. Theorem 2.3 is now fully proved.